Velcro Saddles is contemplating the acquisition of Skiers’ Airbags Inc. The values of the two companies as separate entities are $26 million and $13 million, respectively. Velcro Saddles estimates that by combining the two companies, it will reduce marketing and administrative costs by $530,000 per year in perpetuity. Velcro Saddles is willing to pay $17 million cash for Skiers’. The opportunity cost of capital is 10%.

a. What is the gain from the merger? (Enter your answer in millions rounded to 2 decimal places.)

b. What is the cost of the cash offer? (Enter your answer in millions.)

c. What is the NPV of the acquisition under the cash offer? (Do not round intermediate calculations. Enter your answer in millions rounded to 2 decimal places.)
Answer #1
a Gain from the merger = (26+13+0.53/10%)-(26+13) $ 5.30 Million
b Cost of the cash offer = 17-13 $ 4 Million
c NPV of the acquisition = 5.3-4 $ 1.30 Million